OpenStudy (anonymous):

x-1=√5x-9 solve for x. after you square on both sides what do you do?

OpenStudy (mrnood):

\[x-1 = \sqrt{5x}-9\] OR \[x-1= \sqrt{5x-9}\] ?

OpenStudy (anonymous):

The square root is over 5x-9

OpenStudy (mrnood):

yes it's right IF the original equation was written like the SECOND option I wrote i.e. the 9 is inside the sqrt sign OK So now expand the left hand side (x-1)^2=???

OpenStudy (anonymous):

x^2-2x+1

OpenStudy (mrnood):

good so x^2-2x+1=5x-9 rearrange so it is ??=0

OpenStudy (anonymous):

x^2-7x+8=0 ?

OpenStudy (mrnood):

not quite the last term is incorrect x^2-7x+10=0 Now solve (probably using the quadratic equation formula...) Remenber there are 2 answers to a quadratic

OpenStudy (anonymous):

okay thank you

OpenStudy (mrnood):

actually - it's easier to solve by factorising...

OpenStudy (anonymous):

That was what I was going to do

OpenStudy (mrnood):

Not all quadratics factorise - this one does The formula allows you to solve all solvable quadratics